OpenStudy (anonymous):

which is an equation of the line passing through (3,9) with the slope -2/3?

OpenStudy (anonymous):

\[y=y_1=m(x-x_1)\] \[y-9-=-\frac{2}{3}(x-3)\] \[y-9=-\frac{2}{3}x+6\] \[y=-\frac{2}{3}x+15\]

OpenStudy (anonymous):

point slope formula: \[(y-y_{0}) = m((x-x_{0}))\] plug in the point for \[(x_0, y_0)\] and then solve for y

OpenStudy (anonymous):

Satellite you have a problem in your solution.. \[-\frac{2}{3}(x-3) \ne -\frac{2}{3}x+6\]

OpenStudy (anonymous):

your telling me! it is really \[y=-\frac{2}{3}x+11\]!